If the return of the investment is 8.20% and the inflation rate is 3.12%, what is the inflation-adjusted rate of return?
Answer #1

Inflation Adjusted rate of return

  • It is the return accounts for the effect of inflation on an investment's performance over time.
  • It is also known as the real return

  Inflation Adjusted rate of return

=( 1+ return ) / (1+ inflation rate) -1

Return on the investment =8.20%

Inflation rate = 3.12 %

Infaltion Adjusted rate of return = (1+ 0.082) / (1+0.0312) -1

= 1.082 / 1.0312 -1

=1.049 -1

=0.049 or 4.9%